40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4268     Entry time: Thu Feb 10 05:06:35 2011
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: beat noise : a little bit better, and 1Hz peak from amplitude noise coupling 

 I repeated the same measurement as that Koji did before (see here) with the mixer-based frequency discriminator.

The frequency fluctuation of the beat note is now 50 kHz in rms integrated down to 0.1 Hz, which is a bit better than before.

However there still is the same undesired structure in the spectrum below 10 Hz.

 

 Screen_shot_2011-02-10_at_4.01.43.png

Fig.1 power spectra of the green beat note fluctuation in terms of frequency fluctuation.

   Red curves were taken when the IR was locked to the MC, and the green was locked to the X arm.

Blue curves were taken when both the IR and the green were locked to the X arm.

Black curve was also the one taken when the IR and the green were locked to the X arm, but showing the lower noise level.

I have no idea what exactly was going on when I took the black curve, but this noise level sometimes showed up.

The discrepancy may come from a kind of calibration error although I kept using the same calibration factor to convert the data from count to frequency.

Need more investigations.

 


 Additionally Koji and I took the coherence between the beat fluctuation and the transmitted lights of both the IR and the green.

It showed a strong coherence at 1 Hz, which is one of the dominant noise of the beat note.

This probably indicates that the 1 Hz peak is produced by a coupling from amplitude fluctuation.

Screen_shot_2011-02-10_at_4.52.13.png

 For monitoring the green transmitted light, I used the Jenne's PD (see here)

ELOG V3.1.3-